ganeshie8 (ganeshie8):

Evaluate \(\rm \int\limits_3^5 f^{-1}(x) dx \) if \(\rm f(x) = x^5+x+3\)

OpenStudy (blackops2luvr):

ganeshie... why yew do this?

OpenStudy (yanasidlinskiy):

Umm..^ What if its for school? It's a learning site, isn't it? Or helping site?

OpenStudy (anonymous):

haven't done Calc in a while lol

OpenStudy (ikram002p):

i like that song

OpenStudy (blackops2luvr):

Hes an adult XD

OpenStudy (anonymous):

BOTH

OpenStudy (matlee):

the answer is 5% +6% %&#%$^%^%&#$&^*&*%^7

OpenStudy (catlover5925):

??? @thomaster @Preetha @jordanloveangel

OpenStudy (justnick09):

Thank you @ikram002p

OpenStudy (ikram002p):

hmm interesting ... i cant separate x and y , so whats in ur mind ?

ganeshie8 (ganeshie8):

Somebody claimed this can be done neatly using integration by parts. I have worked using some other convoluted method...

OpenStudy (ikram002p):

so u have an answer hmm

ganeshie8 (ganeshie8):

yes, by using "geometry"

OpenStudy (yanasidlinskiy):

Llalala!!!! @ganeshie8 may we plzzzzzz switch brains?

OpenStudy (ikram002p):

oh , thats ovc but algebraically hmmm

ganeshie8 (ganeshie8):

for sure yes lol im still looking for a solution using IBP or any other method that doesn't require graphing

OpenStudy (yanasidlinskiy):

Yaya!!! Lol, Lets do it today and ummm 2:00pm at the surgery room, haha!!! I'm just kidding. Kk. I'll stop now.

OpenStudy (ikram002p):

u mean IVP ?

OpenStudy (ikram002p):

i think it would work but that would give a series something ... xD

ganeshie8 (ganeshie8):

I have been trying something like below : say \(\large \rm y = f(x) = x^5+x+3\), I need to evaluate \[\large \rm \int_3^5f^{-1}(x)~dx\] substituting \(\rm x = y \) gives me \[\large \rm \int_3^5f^{-1}(f(x))~dy\]

ganeshie8 (ganeshie8):

which is same as \[\large \rm \int_3^5\!x~dy\]

OpenStudy (ikram002p):

haha maybe lets see if it gonna work http://prntscr.com/560gqh

OpenStudy (ikram002p):

brb

ganeshie8 (ganeshie8):

geometrically thats same as : \[5\times 1 - \int_0^1 y~ dx\] \[5\times 1 - \int_0^1 x^5+x+3~~ dx\] this looks a lot like IBP, but i was forced to use geometry for adding/subtracting areas

OpenStudy (anonymous):

\[\begin{matrix} u=f^{-1}(x)&&&dv=dx\\\\ du=\frac{1}{f'\left(f^{-1}(x)\right)}~dx&&&v=x \end{matrix}\]

ganeshie8 (ganeshie8):

do i get \[\rm \int f^{-1}(x) dx = xf^{-1}(x) - \int x \cdot \dfrac{1}{f'\left(f^{-1}(x)\right)} dx\] ?

OpenStudy (anonymous):

Yep, but there's still the matter of not "knowing" what \(f^{-1}(x)\) is. But if we did, we'd be able to figure out the integral from the start.

ganeshie8 (ganeshie8):

does the definite integral help ?

ganeshie8 (ganeshie8):

do i get \[\rm \int_3^5 f^{-1}(x) dx = xf^{-1}(x)\Bigg|_3^5 - \int_3^5 x \cdot \dfrac{1}{f'\left(f^{-1}(x)\right)} dx\] ?

ganeshie8 (ganeshie8):

I can certainly work \(\rm f^{-1}(5)\) and \(\rm f^{-1}(3)\) using f(x), not a problem. Evaluating the rightside integral looks like a pain though

OpenStudy (anonymous):

Yes, \[\int_a^b u~dv=uv\bigg|_a^b-\int_a^b v~du\]

OpenStudy (ikram002p):

what is IBP ?

ganeshie8 (ganeshie8):

integration by parts

OpenStudy (ikram002p):

-.-

OpenStudy (anonymous):

For the second integral, you could consider a substitute \(u=f^{-1}(x)\), so that \(f(u)=x\) and \(f'(u)~du=dx\), if I have the right differentials. Then \[\int_3^5\frac{x}{f'(f^{-1}(x))}~dx=\int_{f(3)}^{f(5)}\frac{f(u)}{f'(f(u))}~f'(u)~du\] if that helps...

OpenStudy (anonymous):

lol.:P IBP.:D

OpenStudy (ikram002p):

its sounds like the same issue of morning problem ! nice

ganeshie8 (ganeshie8):

you meant this rihgt ? XD \[\int_3^5\frac{x}{f'(f^{-1}(x))}~dx=\int_{f^{-1}(3)}^{f^{-1}(5)}\frac{f(u)}{f'(u)}~f'(u)~du \]

OpenStudy (ikram002p):

i think it works good , main idea is to get dx with respect to dy

OpenStudy (anonymous):

Yes that's it, thanks

ganeshie8 (ganeshie8):

Beautiful !!! its IBP all the way xD thanks everyone!

ganeshie8 (ganeshie8):

I think we can generalize

ganeshie8 (ganeshie8):

\[\rm \int_a^b f^{-1}(x) dx = xf^{-1}(x)\Bigg|_a^b - \int_{f^{-1}(a)}^{f^{-1}(b)} \! f(x) dx\]

ganeshie8 (ganeshie8):

i hope it works..

OpenStudy (ikram002p):

but are u sure that 3,5 are valid in this integral xD

ganeshie8 (ganeshie8):

wym by valid ?

OpenStudy (ikram002p):

in the graph u showed me it sound like no shaded area with range of 3,5 :3

OpenStudy (ikram002p):

no no no forget it xD

ganeshie8 (ganeshie8):

here is the graph im using if you want http://prntscr.com/560rnd

ganeshie8 (ganeshie8):

the shaded area is the area we're subtracing - the \(\int vdu\) component in IBP

OpenStudy (ikram002p):

:)

OpenStudy (anonymous):

The general form works for at least a few examples. I tried it with \(f(x)=x^2\) over \([0,2]\). No counter-ex's yet.

myininaya (myininaya):

I think this works \[\int\limits_{3}^{5}f^{-1}(x) dx \\ u=f^{-1}(x) \\ f(u)=x \\ f'(u) du=dx \\ \int\limits_{0}^{1} u f'(u) du \\ \int\limits_{0}^{1} u (5u^4+1) du=\int\limits_{0}^{1}(5u^5+u) du\]

myininaya (myininaya):

I didn't do integration by parts but it seems like a sub works

OpenStudy (anonymous):

@myininaya limits of integration?

ganeshie8 (ganeshie8):

Wow! that looks elegant xD \[\begin{align} \int_a^b f^{-1}(x) dx &= xf^{-1}(x)\Bigg|_a^b - \int_{f^{-1}(a)}^{f^{-1}(b)} \! f(x) dx \\~\\ &=\int_{f^{-1}(a)}^{f^{-1}(b)} \!x f'(x) dx \\~\\ \end{align} \]

myininaya (myininaya):

\[f^{-1}(5)=a \\ f(a)=5 \\ a^3+a+3=5 \\ a=1 \\ f^{-1}(5)=1 \\ f^{-1}(3)=b \\ f(b)=3 \\ b^3+b+3=3 \\ b=0 \\ f^{-1}(3)=0\] that is how I got the limits @Princer_Jones

OpenStudy (anonymous):

great..@myininaya ,.:)